# Hierarchical Neural Simulation-Based Inference Over Event Ensembles

### v1

@@ Abstract
When analyzing real-world data it is common to work with event ensembles, which comprise sets of observations that collectively constrain the parameters of an underlying model of interest. Such models often have a hierarchical structure, where “local” parameters impact individual events and “global” parameters influence the entire dataset. We introduce practical approaches for optimal dataset-wide probabilistic inference in cases where the likelihood is intractable, but simulations can be realized via forward modeling. We construct neural estimators for the likelihood(-ratio) or posterior and show that explicitly accounting for the model’s hierarchical structure can lead to tighter parameter constraints. We ground our discussion using case studies from the physical sciences, focusing on examples from particle physics (particle collider data) and astrophysics (strong gravitational lensing observations).
@@ 1 Introduction
The complexity of scientific models and high-dimensionality of datasets has led to a recent surge in interest in implicitly-specified models, where the likelihood function is intractable but simulations can be realized via mechanistic forward modeling. The paradigm of simulation-based inference (SBI), augmented using tools from machine learning and differentiable optimization more broadly, has emerged as a powerful approach for performing inference in such scenarios [7]. However, the majority of existing methods for simulation-based inference are designed for learning from individual data points and do not fully capitalize on the hierarchical structure of the data-generating process.
To address these gaps, we propose a set of novel approaches that augment existing simulation-based inference techniques, in both frequentist and Bayesian paradigms, with the goal of exploiting the hierarchical structure of the governing models. We contextualize our discussion using case studies from the physical sciences, with a particular focus on particle physics and astrophysics (particle collider data and strong gravitational lensing images, respectively).
Our primary contributions are the following:
  • We substantiate the fact that optimal inference in implicit models with a hierarchical structure requires a dataset-wide approach, contrasted with the more common paradigm of combining implicit likelihood or posterior estimators associated with individual observations;
  • We introduce models for dataset-wide learning that can be used to perform simulation-based inference over event ensembles and can deal with datasets of varying cardinality. We connect our insights to several common use cases in the physical sciences and show how popular simulation-based inference paradigms can be adapted for optimal dataset-wide learning;
@@
Related work: The methods presented herein build upon or complement previous studies across several disciplines. A likelihood-free algorithm for obtaining frequentist profile likelihood ratios corresponding to individual observations was previously presented in Heinrich [13]. Nachman and Thaler [18] explored dataset-wide inference in the context of particle collider studies, without distinguishing nuisance parameters and parameters of interest. Hierarchical inference for implicit models has been studied using Gaussian posterior density estimators in Wagner-Carena et al. [30, 31], where per-event posteriors are learned and combined to obtain estimates for the global parameters. Agrawal and Domke [1] considered amortized variational inference for a simple class of hierarchical models. While similar in spirit, we consider hierarchical inference in the language of modern scientific simulation-based inference methods, and discuss subtleties associated with treating classes of variables in different hierarchies as nuisance parameters. Hierarchical Neural Posterior Estimation [23] is a recently-introduced simulation-based inference method closely related to the deep set-based Bayesian inference model used here. We complement this approach by introducing frequentist analogues for dataset-wide learning; by including amortized estimators for varying dataset length; and by comparing the proposed approach to more traditional methods such as MCMC.
2 Theory and architectures
@@
2.1 Background and theory
Hierarchical models: The joint probability distribution of a set of events with cardinality $N$, with global parameters of interest $\theta$, global nuisance parameters $\theta_{\nu}$ as well as local parameters $z_{i}$ can be written as:
@@ 2.1 Background and theory
Here, $p(N\mid\theta)$ represents the probability of observing $N$ events given the global parameters $\theta$. Often, $p(N\mid\theta)$ is a Poisson distribution, and the full model is known as a marked Poisson process. Our proposed techniques apply to both situations, and we present example applications described by Eqs. (1) as well as (2) below.
Bayesian inference: In the Bayesian paradigm, we introduce a prior $p(\{z\},\theta,\theta_{\nu})=p(\theta,\theta_{\nu})\prod_{i}p(z_{i}\mid\theta,\theta_{\nu})$ and wish to target the posterior distribution over the global as well as local parameters of interest (and marginalizations thereof) given a dataset.
@@
Frequentist inference: In the frequentist setting, inference on parameters is typically split into procedures for point estimation and interval estimation. A popular choice for a point estimator is the maximum-likelihood estimator (MLE), due to its favorable asymptotic property of being the minimum-variance unbiased estimator. Interval estimation and hypothesis testing are based on a subjective choice of a test statistic $t_{\theta,\theta_{\nu}}(x)$ and a desired confidence level:
@@
Simulation-based inference: When the likelihood is not tractable but simulations can be realized via forward modeling, neural simulation-based inference methods can be leveraged to perform inference. In cases where event-level inference quantities can be combined to yield a correct dataset-wide quantity, neural posterior or likelihood-ratio estimators trained on individual events can be combined. For example, marginal posteriors $p(\theta,\theta_{\nu}\mid x_{i})$ can be trained and combined using Eq. (4).
@@
Alternatively, a per-event likelihood ratio estimator parameterized by all parameters, including nuisance parameters, can be learned [6, 18] to train a neural estimate for the per-event likelihood ratio $s_{\phi}(x_{i},\theta,\theta_{\nu})=\frac{p(x_{i}\mid\theta,\theta_{\nu})}{p(x_{i}\mid\theta_{0},\theta_{\nu,0})}$. These can be used to build a dataset-wide likelihood ratio $\prod_{i}\frac{p(x_{i}|\theta,\theta_{\nu})}{p(x_{i}|\theta_{0},\theta_{\nu,0})}$ that can be an input for a dataset-wide statistical analysis to yield posteriors or likelihood ratios which can then be marginalized or profiled over to yield dataset-level quantities. While this simplifies the training procedure as the networks are only performing inference for single events, a disadvantage of this approach is that the downstream statistical treatment can be very costly if the number of nuisance parameters $\theta_{\nu}$ is large, so that the amortization gains from the neural estimates are not fully exploited. In the following we explore simulation-based inference procedures that yield fully amortized dataset-wide inferences.
2.2 Architectures
We consider the following architectures, which respect the hierarchical structure of the forward models and can process event ensembles of varying cardinality.
Deep set-based model:
@@ 2.2 Architectures
Here we consider a simple deep sets-inspired [32] architecture, where per-event encoded features $s^{\mathrm{glob}}_{\phi}(x_{i})$ are aggregated via a permutation-invariant pooling and passed through a decoder network $g_{\psi}$ that outputs the parameters of a variational posterior distribution for the global parameters $q^{\mathrm{glob}}_{\varphi}(\theta\mid\{x\})$. At the same time, local per-event embeddings $s^{\mathrm{loc}}_{\phi}(x_{i})$ are used to condition a density estimator for the local parameters, $q^{\mathrm{loc}}_{\varphi^{\prime}}\left(z\mid\{x\},\theta\right)$, including a conditional dependence on the global parameter $\theta$. In practice, the local and global embeddings are obtained by chunking a feature vector obtained with $s_{\phi}$. For the variational ansatz we consider either a multivariate normal distribution or a normalizing flow [22], with the decoder outputting either the mean-covariance of the normal distribution or the conditioning context for a normalizing flow. The (negative) sum of the local and global posterior log-densities is used as the minimization objective,
@@
For models of the form of Eq. (2) the cardinality of the input set is drawn as $N\sim p(N\mid\theta)$ in the training. The density estimator parameters $\{\varphi,\varphi^{\prime}\}$ as well as the embedding network and decoder parameters $\phi$ and $\psi$ are optimized simultaneously. We note that this is similar to the set-up used in Rodrigues et al. [23].
Transformer-based model:
In a similar vein, we consider a decoder-only transformer [28] which takes feature embeddings from individual events and produces cumulative posterior embeddings using blocks of self-attention and dense layers, applying a causal mask to ensure that output embeddings are only sensitive to preceding events in the sequence. The embeddings are then used to predict the posterior mean and covariance of the variational Gaussian ansatz. The advantage of the transformer approach is that, assuming a uniform distribution on the cardinality $N$, we don’t need to vary the cardinality of the input set during training – since the loss consists of a sum of log-likelihoods corresponding to cumulative posteriors, all cardinalities are considered together; $-\mathcal{L}=\sum_{N}\log q^{\mathrm{glob}}_{\varphi}(\theta\mid\{x\}_{i=1}^{N})+\sum_{i}\log q^{\mathrm{loc}}_{\varphi^{\prime}}(z_{i}\mid x_{i},\theta)$, leaving the data-point encodings implicit.
Fig. 1 shows a schematic illustration of the deep set architecture used in this work. Networks that used LSTM cells to iteratively update a posterior embedding based on incoming data embeddings were also considered, but their convergence properties were noticeably poorer than those of the deep set- and transformer-inspired architectures described above.
@@ 3 Experiments
We describe several case studies of dataset-wide learning, ranging from illustrative “toy” experiments to more prototypical examples representing problems in particle physics and astrophysics. We refer to Appendix A for additional details on the experiments, including details on training. All experiments are implemented using PyTorch [21].
@@ 3.1 Simple multi-variate normal likelihood
The forward model: To verify the ability of our method to recover the true posterior distribution for sets of varying cardinality, we consider a simple multivariate normal likelihood with known covariance matrix $\Sigma$; $p(\{x\}\mid\theta,\Sigma)=\prod_{i=0}^{N}\mathrm{No}(x_{i}\mid\theta,\Sigma)$.
The model parameter is the mean vector $\theta$ , and $\mu_{0},\Sigma_{0}$ are the hyperparameters of the prior multivariate normal distribution. In this case, the posterior distribution $p(\theta\mid\{x\})$ is also a multivariate normal – the prior and posterior are conjugate – with mean $\mu_{\mathrm{post}}$ and covariance $\Sigma_{\mathrm{post}}$ of an updated posterior given by $\mu_{\mathrm{post}}=\left({\Sigma}_{0}^{-1}+N{\Sigma}^{-1}\right)^{-1}\left({\Sigma}_{0}^{-1}{\mu}_{0}+N{\Sigma}^{-1}\overline{{x}}\right)$ and $\Sigma_{\mathrm{post}}=\left({\Sigma}_{0}^{-1}+N{\Sigma}^{-1}\right)^{-1}$ where $\overline{{x}}$ is the sample mean of $x_{i}$ and $N$ is the total cardinality of the dataset.
Inference: We choose $\Sigma=\operatorname{diag}(2,4,6)$, and a data-dimensionality of 5 per parameter dimensions; each data point then consists of 15 features. We train the deep set and transformer architectures described in Sec. 2.2 on 50,000 samples drawn from this likelihood with prior $p(\mu)=\operatorname{No}(0,3)$ and a maximum sequence length of 200. The cardinality of the training set is randomly varying as $N\sim\mathrm{Unif}(1,N_{\mathrm{max}})$. The model is then tested on 500 new sequences, and the distribution (median and middle-68% containment) over inferred standard deviation $\sigma$ for each of the 3 parameters is shown in Fig. 2 (left: deep set, middle: transformer) compared to the true expected scaling of the parameters (dashed lines).
Results: The deep set model typically gives more faithful posterior widths compared to the transformer, potentially due to the simple nature of the problem combined with the more data-hungry nature of the transformer architecture. The right plot shows the evolution of a posterior for a specific sequence using the deep set model, illustrating convergence of the posterior mass around the true point as more data points are included. We note that although this is a simple, analytically tractable example, the final posterior is obtained by analyzing a dataset consisting of 3000 data points – far from a trivial task.

### v2

@@ Abstract
When analyzing real-world data it is common to work with event ensembles, which comprise sets of observations that collectively constrain the parameters of an underlying model of interest. Such models often have a hierarchical structure, where “local” parameters impact individual events and “global” parameters influence the entire dataset. We introduce practical approaches for frequentist and Bayesian dataset-wide probabilistic inference in cases where the likelihood is intractable, but simulations can be realized via a hierarchical forward model. We construct neural estimators for the likelihood(-ratio) or posterior and show that explicitly accounting for the model’s hierarchical structure can lead to significantly tighter parameter constraints. We ground our discussion using case studies from the physical sciences, focusing on examples from particle physics and cosmology.
@@ 1 Introduction
The complexity of scientific models and high-dimensionality of datasets has led to a recent surge in interest in implicitly-specified models, where the likelihood function is intractable but simulations can be realized via mechanistic forward modeling. The paradigm of simulation-based inference (SBI), augmented using tools from machine learning and differentiable optimization more broadly, has emerged as a powerful approach for performing inference in such scenarios (Cranmer et al., 2020). However, the majority of existing methods for simulation-based inference are designed for learning from individual data points and do not fully capitalize on the hierarchical structure of the data-generating process.
To address these gaps, we propose a set of novel approaches that augment existing simulation-based inference techniques, in both frequentist and Bayesian paradigms, with the goal of exploiting the hierarchical structure of the governing models. We contextualize our discussion using case studies from the physical sciences, with a particular focus on particle physics (particle collider data) and astrophysics (strong gravitational lensing images).
In the present work we make several methodological contributions that are necessary for optimal deployment of simulation-based inference methods on large-scale scientific datasets. Our primary contributions are the following:
  • We substantiate theoretically as well as empirically the fact that hierarchy-aware inference in many implicit models with a hierarchical structure requires a dataset-wide approach, contrasted with the more common paradigm of combining implicit likelihood or posterior estimators associated with individual observations. By “hierarchy-aware inference”, we refer to building a still approximate posterior or likelihood-ratio estimator that nevertheless takes into account the full (in our case) hierarchical structure of the assumed forward model. We systematically derive conditions under which such a hierarchical approach is necessary for correct inference, depending on how the model parameters are partitioned into local and global parameters of interest and nuisance parameters;
  • We introduce frequentist as well as Bayesian methods for dataset-wide learning that can be used to perform simulation-based inference over event ensembles and can deal with datasets of varying cardinality. We connect our insights to several common use cases in the physical sciences and show how popular simulation-based inference paradigms can be adapted for optimal dataset-wide learning. In particular, we introduce the first approach for end-to-end frequentist simulation-based inference targeting hierarchical set-valued data;
@@
The paper is organized as follows. In Sec. 2 we provide a background discussion of hierarchical frequentist as well as Bayesian inference, and describe the class of forward models we target. We discuss related work in context of this background Sec. 3. In Sec. 4 we describe the methods and architectures used in this paper. Section 5 presents several case studies including a simple toy example, a particle physics example, and an example from astrophysics, in increasing order of complexity. We conclude in Sec. 6.
2 Theory and background
@@
2.1 Hierarchical models
The joint probability distribution of a set of events with cardinality $N$, with global parameters of interest $\theta$, global nuisance parameters $\theta_{\nu}$ as well as local parameters $z_{i}$ can be written as:
@@ 2.1 Hierarchical models
Here, $p(N\mid\theta)$ represents the probability of observing $N$ events given the global parameters $\theta$. Often, $p(N\mid\theta)$ is a Poisson distribution, and the full model is known as a marked Poisson process. Our proposed techniques apply to both situations, and we present example applications described by equation 1 as well as equation 2 below.
2.2 Bayesian inference
In the Bayesian paradigm, we introduce a prior $p(\{z\},\theta,\theta_{\nu})=p(\theta,\theta_{\nu})\prod_{i}p(z_{i}\mid\theta,\theta_{\nu})$ and wish to target the posterior distribution over the global as well as local parameters of interest (and marginalizations thereof) given a dataset.
@@ 2.2 Bayesian inference
2.3 Frequentist inference
In the frequentist setting, inference on parameters is typically split into procedures for point estimation and interval estimation. A popular choice for a point estimator is the maximum-likelihood estimator (MLE), due to its favorable asymptotic property of being the minimum-variance unbiased estimator. Interval estimation and hypothesis testing are based on a subjective choice of a test statistic $t_{\theta,\theta_{\nu}}(x)$ and a desired confidence level:
@@ 2.4 Simulation-based inference
When the likelihood is not tractable but simulations can be realized via forward modeling, neural simulation-based inference methods can be leveraged to perform inference. In cases where event-level inference quantities can be combined to yield a correct dataset-wide quantity, neural posterior or likelihood-ratio estimators trained on individual events can be combined. For example, marginal posteriors $p(\theta,\theta_{\nu}\mid x_{i})$ can be trained and combined using equation 4.
@@
Alternatively, a per-event likelihood ratio estimator parameterized by all parameters, including nuisance parameters, can be learned (Cranmer et al., 2015; Nachman & Thaler, 2021) to train a neural estimate for the per-event likelihood ratio $s_{\phi}(x_{i},\theta,\theta_{\nu})=\frac{p(x_{i}\mid\theta,\theta_{\nu})}{p(x_{i}\mid\theta_{0},\theta_{\nu,0})}$, These can be used to build a dataset-wide likelihood ratio $\prod_{i}\frac{p(x_{i}|\theta,\theta_{\nu})}{p(x_{i}|\theta_{0},\theta_{\nu,0})}$ that can be an input for a dataset-wide statistical analysis to yield posteriors or likelihood ratios which can then be marginalized or profiled over to yield dataset-level quantities. While this simplifies the training procedure as the networks are only performing inference for single events, a disadvantage of this approach is that the downstream statistical treatment can be very costly if the number of nuisance parameters $\theta_{\nu}$ is large, so that the amortization gains from the neural estimates are not fully exploited.
Given these nuances, methods for set-wide simulation-based inference necessitate architectures which can efficiently model joint posteriors or likelihood ratio over global and local parameters of interest, for sets of varying cardinality, without requiring explicit parameterization over a potentially large number of nuisance parameters (which would otherwise be necessary for hierarchy-aware inference with event-level estimators). In the following, we describe two such architectures for amortized dataset-wide inferences.
3 Related work
The methods presented in this paper build upon previous studies across several disciplines. A likelihood-free algorithm for obtaining frequentist profile likelihood ratios corresponding to individual observations was previously presented in Heinrich (2022). Also in the context of frequentist inference, Nachman & Thaler (2021) explored dataset-wide inference in the context of particle collider studies, without distinguishing nuisance parameters and parameters of interest.
Hierarchical inference for implicit models has been studied using Gaussian posterior density estimators in Wagner-Carena et al. (2021; 2022), where again individual (per-event) posteriors are learned and combined to obtain estimates for the global parameters. Agrawal & Domke (2021) considered amortized variational inference for a simple class of hierarchical models. While similar in spirit, we consider hierarchical inference in the language of modern scientific simulation-based inference methods, and discuss subtleties associated with treating classes of variables in different hierarchies (local and global) as nuisance parameters.
Hierarchical Neural Posterior Estimation (Rodrigues et al., 2021) is a recently-introduced simulation-based inference method closely related to the deep set-based Bayesian inference model used here. We complement this approach by introducing frequentist analogues for dataset-wide learning; by including amortized estimators that can then target datasets of varying sizes; and by comparing the proposed approach to more traditional methods in terms of inference computational time. Finally, Geffner et al. (2022) introduced a method for efficient set-wide learning by combining score estimators targeting a sub-set of events. Our method is complementary, focusing on the unexplored setting of hierarchical models and bridging applications to domain sciences.
4 Methods
Our goal is to target the graphical models described by equation 1 and equation 2 in a simulation-based inference setting. We wish to construct amortized estimators for the likelihood ratio or posterior corresponding to these models such that they (1) are sensitive to the full hierarchical structure of the forward model – i.e., take a dataset-wide view, treating local and global parameters separately, (2) distinguish between nuisance parameters and parameters of interest at various hierarchy levels, and (3) can process datasets of varying cardinality. While previous related works, discussed in Sec. 3, have considered subsets of these requirements, our goal is to target all three with a particular eye towards scientific applications where event ensembles are common. We consider the following architectures, which satisfy these desiderata.
Deep set-based model
@@ 4 Methods
Here we consider a simple deep sets-inspired (Zaheer et al., 2017) architecture, where per-event encoded features $s^{\mathrm{glob}}_{\phi}(x_{i})$ are aggregated via a permutation-invariant pooling and passed through a decoder network $g_{\psi}$ that outputs the parameters of a variational posterior distribution for the global parameters $q^{\mathrm{glob}}_{\varphi}(\theta\mid\{x\})$ (Papamakarios & Murray, 2016). At the same time, local per-event embeddings $s^{\mathrm{loc}}_{\phi}(x_{i})$ are used to condition a density estimator for the local parameters, $q^{\mathrm{loc}}_{\varphi^{\prime}}\left(z\mid\{x\},\theta\right)$, including a conditional dependence on the global parameter $\theta$. In practice, the local and global embeddings are obtained by chunking a feature vector obtained with $s_{\phi}$. For the variational ansatz we consider either a multivariate normal distribution or a normalizing flow (Rezende & Mohamed, 2015), with the decoder outputting either the mean-covariance of the normal distribution or the conditioning context for a normalizing flow. The (negative) sum of the local and global posterior log-densities is used as the minimization objective,
@@
We derive and further motivate this loss function in App. B. For models of the form of equation 2, the cardinality of the input set is drawn as $N\sim p(N\mid\theta)$ in the training. The density estimator parameters $\{\varphi,\varphi^{\prime}\}$ as well as the embedding network and decoder parameters $\phi$ and $\psi$ are optimized simultaneously. We note that this is similar to the set-up used in Rodrigues et al. (2021). Fig. 1 shows a schematic illustration of this deep set-based architecture.
Transformer-based model
As an alternative to deep set-based aggregation with varying cardinality at training time, we consider a decoder-only transformer (Vaswani et al., 2017; Phuong & Hutter, 2022) which takes feature embeddings from individual events and produces cumulative posterior embeddings using blocks of self-attention and dense layers, applying a causal mask to ensure that output embeddings are only sensitive to preceding events in the sequence. The embeddings are then used to predict the posterior mean and covariance of the variational Gaussian ansatz. A potential advantage of the transformer approach is that, assuming a uniform distribution on the cardinality $N$, we do not need to vary the cardinality of the input set during training. Since the loss consists of a sum of log-likelihoods corresponding to cumulative posteriors, one for each cardinality $N$ up to some specified maximum value, all cardinalities are considered together; $-\mathcal{L}=\sum_{N}\log q^{\mathrm{glob}}_{\varphi}(\theta\mid\{x\}_{i=1}^{N})+\sum_{i}\log q^{\mathrm{loc}}_{\varphi^{\prime}}(z_{i}\mid x_{i},\theta)$, leaving the data-point encodings implicit.
Networks that used LSTM cells to iteratively update a posterior embedding based on incoming data embeddings were also considered, but their convergence properties were noticeably poorer than those of the deep set- and transformer-inspired architectures described above.
@@ 5 Experiments
We describe several case studies of dataset-wide learning, ranging from illustrative “toy” experiments to more prototypical examples representing problems in particle physics and astrophysics. We refer to App. A for additional details on the experiments, including details on training. All experiments are implemented using PyTorch (Paszke et al., 2019).
@@ 5.1 Simple multi-variate normal likelihood
The forward model
To verify the ability of our method to recover the true posterior distribution for sets of varying cardinality, we consider a simple multivariate normal likelihood with known covariance matrix $\Sigma$; $p(\{x\}\mid\theta,\Sigma)=\prod_{i=0}^{N}\mathrm{No}(x_{i}\mid\theta,\Sigma)$.
The model parameter is the mean vector $\theta$, and $\mu_{0},\Sigma_{0}$ are the hyperparameters of the prior multivariate normal distribution; $p(\theta)=\mathrm{No}(\mu_{0},\Sigma_{0})$. In this case, the posterior distribution $p(\theta\mid\{x\})$ is also a multivariate normal – the prior and posterior are conjugate – with mean $\mu_{\mathrm{post}}$ and covariance $\Sigma_{\mathrm{post}}$ of an updated posterior given by $\mu_{\mathrm{post}}=\left({\Sigma}_{0}^{-1}+N{\Sigma}^{-1}\right)^{-1}\left({\Sigma}_{0}^{-1}{\mu}_{0}+N{\Sigma}^{-1}\overline{{x}}\right)$ and $\Sigma_{\mathrm{post}}=\left({\Sigma}_{0}^{-1}+N{\Sigma}^{-1}\right)^{-1}$ where $\overline{{x}}$ is the sample mean of $x_{i}$ and $N$ is the total cardinality of the dataset. We choose $\Sigma=\operatorname{diag}(2,4,6)$, with each sample consisting of 5 draws from the multi-variate normal distribution; each individual data point then consists of 15 features.
Inference
We train the deep set and transformer architectures described in Sec. 4 on 50,000 samples drawn from this likelihood with prior $p(\mu)=\operatorname{No}(0,3)$ and a maximum sequence length of $N_{\mathrm{max}}=200$. The cardinality of the training set is randomly varying as $N\sim\mathrm{Unif}(1,N_{\mathrm{max}})$. Further details on the training procedure are provided in App. A.
Results
The model is then tested on 500 new sequences, and the distribution (median and middle-68% containment) over inferred standard deviation $\sigma$ for each of the 3 parameters is shown in Fig. 2 (left: deep set, middle: transformer) compared to the true expected scaling of the parameters (dashed lines). We show $\sigma_{1}$, $\sigma_{2}$, and $\sigma_{3}$, which are the diagonal entries of $\Sigma_{\mathrm{post}}$, the covariance of the posterior on the target mean parameters $\theta$. The deep set model typically gives more faithful posterior widths compared to the transformer, potentially due to the relatively simple nature of the problem combined with the more data-hungry nature of the transformer architecture. Given this, we restrict our experiments in subsequent sections to use the deep set-based architecture, noting that the transformer approach may nevertheless yield advantages depending on the specific structure of the forward model. The right plot shows the evolution of a posterior for a specific sequence using the deep set model, illustrating convergence of the posterior mass around the true point as more data points are included. We note that although this is a simple, analytically tractable example, the final posterior is obtained by analyzing a dataset of dimensionality 3000 – far from a trivial task.